Walter Company issues $750,000 of 12% bonds that pay interest semiannually and mature in 10 years. Compute the bonds’ issue price assuming that the bonds’ market interest rate is:

  1. 10% per year compounded semiannually
  2.    14% per year compounded semiannually
Answer #1

Answer:

A.
10% per year compounded semiannually
Semiannual interest rate = 12/2 = 6%
Semiannual market rate = 10/2 = 5%
Number of semiannual period = 10*2 = 20
Issue price
= Semiannual interest*Sum of PV factor @5% for 20 period + Redemption value*PV factor @5% for 20 period
= (750000*0.06)*12.462 + 750000*0.377
= $843,540
b.      14% per year compounded semiannually
Semiannual interest rate = 12/2 = 6%
Semiannual market rate = 14/2 = 7%
Number of semiannual period = 10*2 = 20
Issue price
= Semiannual interest*Sum of PV factor @7% for 20 period + Redemption value*PV factor @7% for 20 period
= (750000*0.06)*10.594 + 750000*0.258
= $670,230